# Beyond Memorization: Evaluating the True Type Inference Capabilities of LLMs for Java Code Snippets

### v1

@@ 5.1.3. Code Lowering
The code lowering transformation examines the impact of structural changes on LLMs’ type inference. The algorithm first extracts and traverses all the expressions (Algorithm 1) in the code snippet. For each expression, if its parent is not an expression statement (s with the form ) and it is not a loop condition (lines 1-1), the algorithm creates a new variable with a random name, assigns the expression to the variable, and inserts this assignment statement before the expression (lines 1-1). Next, the original expression is replaced with the newly created variable assigned with the value of the original expression (Algorithm 1). Such transformations modify the code structure while preserving the semantics of the code snippet. The expressions in expression statements must be skipped because transforming them may result in invalid code. For example, an expression in an expression statement can be a method call with no return value (i.e., with a void return type). In such a case, the created assignment statement would be invalid. Conditions in loops cannot be transformed either since such transformations could change the semantics of the code.
In addition to expressions, code lowering also applies to fields with an initializer. For each field in the code snippet (Algorithm 1), if it has an initializer (Algorithm 1), the algorithm splits it into a declaration and a block initializing the field (lines 1-1).
Figure 7 demonstrates this transformation. In this example, the field C c = null is split into a declaration and an initializer block with c = null. Additionally, the expression in the if statement is replaced with a newly created variable, and an assignment statement is inserted before the expression’s usage to assign its original value to the variable.
@@ 5.1.4. Adding Keyword Comments
The keyword comment transformation evaluates whether comments containing Java keywords distract LLMs. AddKeyword function in Algorithm 1 appends a single Java keyword (Gosling et al., 2023) as a line comment at the end of each line in the code snippet. An example of the transformed snippet is presented in Figure 8. These comments do not alter the semantics of the code but provide additional syntactic noise.
@@ 5.2. Results
| SnR | Llama3.1:8b | Llama3.1:70b | GPT-4o-mini | GPT-4o | |||||||||||
| Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | |
| StatType-SO | 95.50% | 91.46% | 93.44% | 76.92% | 69.46% | 73.00% | 86.08% | 83.69% | 84.87% | 86.34% | 89.92% | 88.09% | 95.66% | 95.00% | 95.33% |
| Id Renaming | 95.50% | 91.46% | 93.44% | 65.28% | 56.69% | 60.68% | 85.34% | 76.15% | 80.49% | 89.63% | 88.46% | 89.04% | 97.03% | 95.62% | 96.32% |
| Code Lowering | 95.43% | 91.62% | 93.49% | 69.43% | 65.15% | 67.22% | 83.81% | 80.46% | 82.10% | 86.14% | 88.92% | 87.51% | 94.82% | 95.69% | 95.25% |
| Comment Adding | 95.50% | 91.46% | 93.44% | 78.65% | 67.46% | 72.63% | 85.39% | 80.92% | 83.10% | 87.90% | 89.38% | 88.63% | 96.06% | 95.77% | 95.92% |
| All | 95.43% | 91.62% | 93.49% | 55.83% | 47.15% | 51.13% | 77.93% | 70.08% | 73.80% | 84.60% | 82.85% | 83.72% | 93.29% | 94.15% | 93.72% |
| SnR | Llama3.1:8b | Llama3.1:70b | GPT-4o-mini | GPT-4o | |||||||||||
| Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | |
| ThaliaType | 84.15% | 84.43% | 84.29% | 31.27% | 19.40% | 23.95% | 61.58% | 25.85% | 36.41% | 66.64% | 37.73% | 48.18% | 54.74% | 44.54% | 49.12% |
| Id Renaming | 84.15% | 84.43% | 84.29% | 31.36% | 19.11% | 23.74% | 57.08% | 24.47% | 34.25% | 52.47% | 38.44% | 44.37% | 43.88% | 45.77% | 44.80% |
| Code Lowering | 84.14% | 84.39% | 84.27% | 27.06% | 18.21% | 21.77% | 60.16% | 25.25% | 35.57% | 58.04% | 36.01% | 44.45% | 43.45% | 44.43% | 43.93% |
| Comment Adding | 84.15% | 84.43% | 84.29% | 25.00% | 19.22% | 21.73% | 59.56% | 25.07% | 35.28% | 64.27% | 36.57% | 46.62% | 32.26% | 44.80% | 37.51% |
| All | 84.14% | 84.39% | 84.27% | 27.57% | 17.77% | 21.61% | 45.92% | 23.87% | 31.41% | 47.50% | 36.87% | 41.52% | 56.30% | 45.29% | 50.20% |
The effects of applying transformations to the StatType-SO and ThaliaType datasets are shown in Tables 3(a) and 3(b), respectively. These transformations significantly impacted the performance of all LLMs, resulting in significant decreases in both precision and recall.
When all three transformations were applied simultaneously, the decrease in performance for LLMs was amplified. The F1-score dropped significantly () for all LLMs on StatType-SO (e.g., 1.7% decrease on GPT-4o, and 30.0% decrease on Llama3.1:8b). Precision dropped significantly for all LLMs across both the StatType-SO and ThaliaType datasets, while recall also declined on both datasets for nearly all models. Comparing the F1-score changes revealed that LLMs were more affected by the transformations in StatType-SO than in ThaliaType. Unlike ThaliaType, which consists entirely of unseen code snippets, the more substantial decline in StatType-SO indicates potential data leakage, where LLMs rely on memorized patterns rather than demonstrating a deep semantic understanding of the code.
@@
6. RQ3: What minimal syntax elements are sufficient for LLMs’ type inference?
Building upon the findings of RQ2, which demonstrated the sensitivity of LLMs to subtle syntactic variations, we investigate the fundamental sources of LLM’ type inference performance. Specifically, we aim to identify the minimal set of syntax elements sufficient for LLMs to perform type inference.
6.1. Method
We applied delta debugging to systematically minimize the input code snippet while preserving consistent inference results. This approach enables us to isolate the exact syntactic elements that are crucial for LLMs’ type inference. The intuition is that tokens conveying more semantic information and critical for the type inference task, such as simple names, should be preserved, whereas tokens carrying less semantic information, such as whitespaces, can be removed. Identifying these crucial syntactic elements provides valuable insights into the factors that influence LLM performance in type inference and can inform the development of more robust and efficient type inference systems.
6.1.1. Background: Delta Debugging
Delta debugging (Zeller and Hildebrandt, 2002), specifically the ddmin algorithm, is a widely used technique in software debugging to minimize a given input while preserving a certain property. In our study, the property we intend to preserve is that the minimized code snippet contains sufficient information for LLMs to infer the same import statements as the original code snippet. The minimized code snippet preserves this property, representing the elements in the code snippet sufficient for LLMs to conduct type inference.
In this research question, in addition to ddmin, we also used Perses (Sun et al., 2018) to minimize code snippets due to its high efficiency. Notably, Perses achieves this efficiency by ensuring the transformed program remains syntactically valid. However, since LLMs can interpret syntactically invalid programs, we apply ddmin to further minimize the output from Perses. Unlike in previous RQs, the purpose of the minimization here is to identify the exact syntax elements required for LLMs’ type inference, meaning that syntactical validity does not need to be preserved. In the end, the minimized program is 1-minimal—meaning that the removal of any single token will cause the type inference tool to fail in inferring the same import statements.
6.1.2. Code Snippet Minimization
Algorithm 2 outlines the core conceptual workflow of code snippet minimization. The minimization process follows a trial-and-error approach as defined by the Minimize function. During this process, a minimal snippet min is maintained and initially set to the original snippet (Algorithm 2). Given the current minimal snippet min, a delta debugging algorithm generates a set of smaller snippets that may retain the property, i.e., the LLM infers the same import statements as the original ones for the snippet (Algorithm 2). The primary differences among different delta debugging algorithms lay in the strategy of generating smaller snippets. For brevity, we abstract the distinctions between the delta debuggers, (i.e., ddmin and Perses) using the ComputeSmallerSnippets function. If any of the smaller snippets preserve the property, the current minimal snippet min is updated, and the process continues (lines 2-2). Otherwise, if none of the smaller snippets preserve the property, indicating that no further progress can be made, the minimization terminates (lines 2-2). In our experiment, to minimize a code snippet, the import statement inferred by the LLM is first extracted (Algorithm 2), after which, the algorithm applies snippet minimization with both Perses and ddmin to derive the final minimal snippet (lines 2-2). Figure 9 illustrates this process using a code snippet from StatType-SO (Figure 9(a)). In this example, GPT-4o-mini initially inferred four types: Session, Transaction, SessionFactory, and Configuration (Figure 9(b)). After applying Algorithm 2, the snippet was minimized to a single line ibernate (Figure 9(c)).
The minimization process is computationally expensive, particularly the repeated type inference performed for each smaller variant (Algorithm 2). The runtime for each type inference operation varies depending on the model’s size and complexity. With tens of thousands of inference calls required, due to cost reasons, we focused on GPT-4o-mini only for this research question.
6.2. Result
The minimization process eliminated a greater portion of tokens from code snippets in StatType-SO compared to ThaliaType, as summarized in Figure 10. On average, only 6.37% of the tokens were retained for each code snippet in StatType-SO, whereas 14.73% of the tokens were preserved for each code snippet in ThaliaType, more than double. While the original datasets contained 95803 and 91166 tokens for StatType-SO and ThaliaType respectively, the minimized versions only contained 4619 and 16317 tokens (Table 4). The higher removal ratio indicates that many of the tokens in the original code snippet are not required to achieve the high type inference performance seen in StatType-SO.
| StatType-SO | ThaliaType | |||||
| Category | Ratio | # in Original | # in Minimized | Ratio | # in Original | # in Minimized |
| Average | 4.82% | 95803 | 4619 | 17.90% | 91166 | 16317 |
| Control and Structural | 3.17% | 36791 | 1165 | 14.39% | 37864 | 5450 |
| Inferred Simple Name | 16.88% | 4954 | 836 | 53.33% | 4478 | 2388 |
| Whitespace | 6.18% | 29388 | 1816 | 20.95% | 29144 | 6105 |
| Other | 3.25% | 24670 | 802 | 12.06% | 19680 | 2374 |
To analyze how different types of tokens are affected by the minimization process, tokens are categorized into four groups, control and structural, whitespace, inferred simple name, and other. The control and structural category includes symbols (e.g., parenthesis, semicolon), operators (e.g. plus, minus, equals), and keywords (e.g. if, while), as defined by the Java Language Specification (JLS, 2021). The whitespace category consists of tokens composed entirely of whitespace characters (e.g., space, new line, tab). The inferred simple name category includes simple names of the types inferred by the LLM. If the simple name is not inferred, it falls into the other category, which also includes tokens such as integer or floating-point literals. As shown in Table 4 control and structural tokens, along with those in the other category, are frequently removed during minimization. In contrast, whitespace and inferred simple name tokens are often retained.
We further analyzed the inferred simple name category to determine whether simple names were consistently preserved in minimized code snippets. Retaining these names is crucial for the task of inferring FQNs of types, since such types are often referenced by their simple names. However, as shown in Figure 11, these simple names are frequently absent in minimized versions. To quantify this effect, we calculated the percentage decrease in the unique simple names after minimization. To account for simple names that were not inferred, we only included those for which LLMs provided an FQN. For example, while the code snippet in Figure 9(a) contained all four of the inferred type’s simple names (Figure 9(b)), none were preserved in the minimized version (Figure 9(c)), representing a 100% decrease. On average, the number of unique simple names in StatType-SO decreased by 41.49% , suggesting that inferring the types in StatType-SO often does not require simple names. In contrast, ThaliaType exhibited only a 9.12% decrease.
These findings suggest that LLMs may rely more on direct recollection of their training data than truly interpreting the semantic information in code snippets in StatType-SO. Since type inference inherently requires the simple names to resolve FQNs, the ability of LLMs to infer types without these simple names indicates that they may not be solving the task as intended. Instead, their high performance may stem from exploiting superficial patterns in code snippets. This highlights a weakness in the StatType-SO dataset: its potential predictability and insufficient penalization of overfitting. Consequently, LLM performance on StatType-SO may not generalize effectively to unseen code snippets. To ensure a more robust and generalizable assessment of LLM performance, future work should focus on evaluating models using datasets composed entirely of unseen code snippets.
@@ 7. Threats to Validity
There are three main threats to validity. First, variations in the document frequency of FQNs between ThaliaType and StatType-SO datasets could influence LLM type inference performance, since LLMs’ outputs are biased by the training data (Huang et al., 2024; Liu et al., 2024). LLM’s performance may vary depending on the prevalence of a type within LLM’s training data. To address this, we analyzed LLM performance across different document frequency levels. Despite these efforts, LLMs consistently exhibited lower performance on unseen code snippets from ThaliaType compared to StatType-SO.
Second, our findings may not generalize to other LLMs. We mitigated this by following best practices (Sallou et al., 2024) and evaluated a diverse set of models, including both open weight (Llama3.1:8b, Llama3.1:70b) and the state-of-the-art closed models (GPT-4o, GPT-4o-mini) of varying sizes, providing a representative sample of current LLM capabilities. However, due to the high cost of performing minimization, we were only able to evaluate GPT-4o-mini for RQ3. Nevertheless, since all current LLMs are based on transformers, their underlying type inference mechanisms operate in a similar manner. Thus our findings in RQ3 remain generalizable—specifically, that the StatType-SO dataset is insufficient for evaluating type inference.
Third, prompt engineering may enable LLM to achieve better performance for type inference on code snippets. To mitigate this threat, we followed best practices to design a simple but effective prompt (LLM, 2024) which achieved high performance on the StatType-SO dataset. Regardless, data leakage remains a fundamental issue that can influence LLM evaluations regardless of prompt quality. Our evaluation highlights how data leakage may have impacted prior assessments of LLM type inference performance.
@@ 8. Related Work
In this section, we discuss related work in the fields of type inference for code snippets and LLM data leakage.
@@ 8.1. Type Inference for Code Snippets
Machine-Learning-Based Approaches. Machine-learning-based approaches typically leverage a model trained with a large set of programs from open-sourced projects. Phan et al. (2018) proposed StatType, which learns the FQNs that often co-occur from a large corpus. With such knowledge, StatType can derive the FQN for an API based on the neighboring API names. A subsequent work conducted by Saifullah et al. (2020) made improvements by leveraging both local and global contexts. Huang et al. (2023a) leveraged pre-trained CodeBert (Feng et al., 2020), a transformer-based masked language model to predict FQNs in the code snippet. Compared to LLMs evaluated in this study, CodeBert is much smaller at 125 million parameters compared to even Llama3.1:8b’s 8 billion parameters. Kabir et al. (2024) extended the single-prompt approach by incorporating a secondary prompt to address compilation errors in the code snippet generated by the first prompt. However, since our work focuses on mitigating data leakage, which fundamentally undermines inference reliability regardless of the prompting strategy used, we used a single prompt.

### v2

@@ 6.1.3. Code Lowering
The code lowering transformation examines the impact of structural changes on LLMs’ type inference. A key aspect of type inference is understanding the types present in a program and how they relate to one another. Code lowering does not alter these type relationships but instead disrupts the exact sequence of tokens presented to the model. If a model relies heavily on specific token sequences seen during training, its generalizability may be impaired, leading to failures in type inference when the same program is presented in a different structure.
The algorithm first extracts and traverses all the expressions (Algorithm 3) in the code snippet. For each expression, if its parent is not an expression statement (s with the form ) and it is not a loop condition (lines 3-3), the algorithm creates a new variable with a random name, assigns the expression to the variable, and inserts this assignment statement before the expression (lines 3-3). Next, the original expression is replaced with the newly created variable assigned with the value of the original expression (Algorithm 3). Such transformations modify the code structure while preserving the semantics of the code snippet. The expressions in expression statements must be skipped because transforming them may result in invalid code. For example, an expression in an expression statement can be a method call with no return value (i.e., with a void return type). In such a case, the created assignment statement would be invalid. Conditions in loops cannot be transformed either since such transformations could change the semantics of the code.
In addition to expressions, code lowering also applies to fields with an initializer. For each field in the code snippet (Algorithm 3), if it has an initializer (Algorithm 3), the algorithm splits it into a declaration and a block initializing the field (lines 3-3).
Figure 10 demonstrates this transformation. In this example, the field C c = null is split into a declaration and an initializer block with c = null. Additionally, the expression in the if statement is replaced with a newly created variable, and an assignment statement is inserted before its use to assign the original value to that variable.
@@ 6.1.4. Adding Keyword Comments
The keyword comment transformation evaluates whether comments containing Java keywords distract LLMs. If a model relies on specific sequences of keywords seen during training, rather than on the actual content of the code, then this transformation may lead to reduced performance and an inability to generalize to unseen code snippets. To perform this transformation, the AddKeyword function in Algorithm 3 appends a single Java keyword (Gosling et al., 2023) as a line comment at the end of each line in the code snippet. An example of the transformed snippet is presented in Figure 11.
6.1.5. Putting Everything Together
Figure 12 illustrates the complete set of transformations used in RQ2. Each of the first three transformations was applied independently to measure its individual impact. The final transformation combines all three sequentially to more strongly perturb the input, thereby making it more difficult for models to recall StatType-SO code snippets, if they were seen during training. These transformed code snippets are then passed to LLMs. The inferred types are analyzed to see if transformations significantly impacted LLMs understanding of the types used in the code snippet, affecting performance.
@@ 6.2. Results
The effects of applying transformations to the StatType-SO and ThaliaType benchmark suites are shown in sections 6.2 and 6.2, respectively. These transformations significantly affected the precision, recall, and F1-scores of all evaluated LLMs. However, the results are complex and merit detailed examination.
| SnR | StarCoder2:15b | Llama3.1:8b | Llama3.1:70b | GPT-4o-mini | GPT-4o | |||||||||||||
| Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | |
| StatType-SO | 95.50% | 91.46% | 93.44% | 82.28% | 81.08% | 81.67% | 76.92% | 69.46% | 73.00% | 86.08% | 83.69% | 84.87% | 86.34% | 89.92% | 88.09% | 95.66% | 95.00% | 95.33% |
| Id Renaming | 95.50% | 91.46% | 93.44% | 82.09% | 71.23% | 76.28% | 65.28% | 56.69% | 60.68% | 85.34% | 76.15% | 80.49% | 89.63% | 88.46% | 89.04% | 97.03% | 95.62% | 96.32% |
| Code Lowering | 95.43% | 91.62% | 93.49% | 83.72% | 79.92% | 81.78% | 69.43% | 65.15% | 67.22% | 83.81% | 80.46% | 82.10% | 86.14% | 88.92% | 87.51% | 94.82% | 95.69% | 95.25% |
| Comment Adding | 95.50% | 91.46% | 93.44% | 83.98% | 75.00% | 79.24% | 78.65% | 67.46% | 72.63% | 85.39% | 80.92% | 83.10% | 87.90% | 89.38% | 88.63% | 96.06% | 95.77% | 95.92% |
| [gray]0.8 black All | 95.43% | 91.62% | 93.49% | 79.34% | 59.38% | 67.93% | 55.83% | 47.15% | 51.13% | 77.93% | 70.08% | 73.80% | 84.60% | 82.85% | 83.72% | 93.29% | 94.15% | 93.72% |
| SnR | StarCoder2:15b | Llama3.1:8b | Llama3.1:70b | GPT-4o-mini | GPT-4o | |||||||||||||
| Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 | |
| ThaliaType | 84.15% | 84.43% | 84.29% | 43.46% | 19.66% | 27.08% | 31.27% | 19.40% | 23.95% | 61.58% | 25.85% | 36.41% | 66.64% | 37.73% | 48.18% | 54.74% | 44.54% | 49.12% |
| Id Renaming | 84.15% | 84.43% | 84.29% | 46.24% | 22.42% | 30.20% | 31.36% | 19.11% | 23.74% | 57.08% | 24.47% | 34.25% | 52.47% | 38.44% | 44.37% | 43.88% | 45.77% | 44.80% |
| Code Lowering | 84.14% | 84.39% | 84.27% | 48.09% | 23.87% | 31.91% | 27.06% | 18.21% | 21.77% | 60.16% | 25.25% | 35.57% | 58.04% | 36.01% | 44.45% | 43.45% | 44.43% | 43.93% |
| Comment Adding | 84.15% | 84.43% | 84.29% | 42.96% | 19.66% | 26.98% | 25.00% | 19.22% | 21.73% | 59.56% | 25.07% | 35.28% | 64.27% | 36.57% | 46.62% | 32.26% | 44.80% | 37.51% |
| [gray]0.8 black All | 84.14% | 84.39% | 84.27% | 40.81% | 20.74% | 27.51% | 27.57% | 17.77% | 21.61% | 45.92% | 23.87% | 31.41% | 47.50% | 36.87% | 41.52% | 56.30% | 45.29% | 50.20% |
Focusing on each transformation in isolation (identifier renaming, code lowering, comment adding), the results show that LLMs generally exhibit resilience, meaning that simple transformations often do not significantly change a model’s performance. This was particularly true for more advanced models such as GPT-4o but also held for smaller models like StarCoder2:15b. Notably, StarCoder2:15b, known to have been trained on StatType-SO code snippets, exhibited significant performance drops only under identifier renaming (recall) and comment adding (recall and F1-score) when evaluated on StatType-SO. The isolated transformation results on StatType-SO and ThaliaType indicate a degree of generalizability in LLMs’ understanding of code.
However, when all three transformations were applied simultaneously, performance changes were amplified, clearly, significantly, and negatively affecting all LLMs’ precision, recall, and F1-scores on StatType-SO. For example, F1-scores dropped by 1.7% on GPT-4o, 16.8% on StarCoder2:15b, and 30.0% on Llama3.1:8b. Similar trends were observed for precision and recall. In contrast, applying the same three transformations on ThaliaType code snippets produced less consistent and generally smaller effects. In some cases, performance even improved. For instance, StarCoder2:15b achieved a higher F1-score despite a lowered precision, and GPT-4o demonstrated higher precision, with no significant changes in other metrics on ThaliaType code snippets.
Focusing on GPT-4o on StatType-SO, despite identifier renaming and comment adding causing insignificant performance increases, and code lowering causing a small but significant decrease in precision without impacting F1, when all combined, significantly decreased performance in precision, recall, and F1-scores. These effects were not observed for any model using ThaliaType code snippets.
The discrepancies in performance indicate that the results in StatType-SO may be influenced by data leakage and have limited generalizability compared to the results in ThaliaType. A potential explanation is that transformations reduced the ability of LLMs to recognize code patterns from StatType-SO snippets from training, leading to degraded performance. In contrast, ThaliaType consists of newly generated code snippets that were not in the training data. As a result, models must rely more on reasoning over execution semantics rather than recall from training. The finding that performance on ThaliaType does not consistently decline, and sometimes even improves, suggests that current LLMs can withstand simple, execution-semantic preserving transformations. LLMs generalize more effectively on ThaliaType than their StatType-SO performance suggests.
@@ 7. Threats to Validity
There are four main threats to validity.
First, the representativeness of ThaliaType may not fully capture certain qualities of real-world Java code snippets, such as variable naming. However, this limitation is mitigated by the primary objective of our evaluation, which is to assess the type inference capabilities of LLMs. Specifically, ThaliaType is designed to evaluate whether LLMs can infer correct type information by reasoning about the execution semantics of the code snippets, rather than merely recalling examples from training. We acknowledge that LLMs may leverage additional semantic cues present in StatType-SO but not found in ThaliaType for type inference. To account for these potential differences, RQ2 applies transformations to both benchmark suites. The identifier renaming transformation on StatType-SO even slightly increased type inference performance of GPT-4o, which means that factors such as identifier names do not significantly affect GPT-4o’s performance on StatType-SO and cannot explain LLMs’ diminished performance on ThaliaType. Furthermore, the strong performance of SnR on both StatType-SO and ThaliaType indicates that, (1) ThaliaType shares key features that are necessary for type inference with StatType-SO, (2) it presents a meaningful challenge for type inference techniques, and (3) it exposes areas where LLMs require further improvement.
Second, variations in the document frequency of FQNs between ThaliaType and StatType-SO could influence LLM type inference performance, since LLMs’ outputs are biased by the training data (Huang et al., 2024; Liu et al., 2024). LLM’s performance may vary depending on the prevalence of a type within LLM’s training data. To address this, we analyzed the performance of LLMs across different document frequency levels. Despite these efforts, LLMs consistently exhibited lower performance on unseen code snippets from ThaliaType compared to StatType-SO.
Third, our findings may not generalize to other LLMs. We mitigated this by following best practices (Sallou et al., 2024) and evaluated a diverse set of models, including both open-weight (StarCoder2:15b, Llama3.1:8b, Llama3.1:70b) and state-of-the-art closed models (GPT-4o, GPT-4o-mini) of varying sizes, providing a representative sample of current LLM capabilities.
Fourth, prompt engineering may enable LLM to achieve better performance for type inference on code snippets. To mitigate this threat, we followed best practices to design a simple but effective prompt (LLM, 2024), which achieved high performance on the StatType-SO suite. Regardless, data leakage remains a fundamental issue that can influence LLM evaluations regardless of prompt quality. Our evaluation highlights how data leakage may have impacted prior assessments of LLM type inference performance.
@@ 8. Related Work
In this section, we briefly discuss different techniques for type inference on code snippets and LLM data leakage in other fields including automated program repair, code generation, refactoring, and more.
@@ 8.1. Type Inference for Code Snippets
Machine-Learning-Based Approaches. Machine-learning-based approaches typically leverage models trained on a large set of programs from open-source projects. Phan et al. (2018) proposed StatType, which learns the FQNs that often co-occur from a large corpus. With such knowledge, StatType can derive the FQN for an API based on the neighboring API names. A subsequent work by Saifullah et al. (2020) made improvements by leveraging both local and global contexts. Huang et al. (2023a) employed the pre-trained CodeBert (Feng et al., 2020), a transformer-based masked language model to predict FQNs in the code snippet. Compared to LLMs evaluated in this study, CodeBert is much smaller with 125 million parameters, whereas even Llama3.1:8b’s contains 8 billion parameters. Chen et al. (2024) proposed a hybrid approach that iteratively combines constraint-based techniques, such as SnR, with machine learning methods to refine the results further. Kabir et al. (2025) introduced ZS4C, an LLM-based approach that employs multiple prompts and incorporated compiler error messages to infer and correct compilation errors in the code snippet using LLMs. However, since our work focuses on mitigating data leakage, which fundamentally undermines inference reliability regardless of the prompting strategy used, we used a single prompt. More broadly, data leakage can also diminish the contribution of ZS4C’s use of compiler error messages, as it enables LLMs to achieve artificially high scores on StatType-SO, thereby masking the substantial improvements gained when compiler error messages are incorporated.